A train is travelling on a uniform speed of 60km/h if the speed was reduced by 10km/h the time for the journey is increased ny 1/2 hour. Find the speed of the train.
1
Expert's answer
2018-08-15T09:05:08-0400
According to the Question the speed was reduced by 10 km/h. Hence

v=60-10=50 ( km)/h.